# Tactile Perception in Upper Limb Prostheses:

### v1

@@ V Tactile Feedback Recognition Experiment
V-A Datasets
New experiments are carried out to obtain four datasets, one for each robotic hand. The prosthetic hands with the reference socket are set by the same supporting structure in Section III and IV. The five IMUs are kept on the inner surface of the socket to measure the vibration pattern after each contact. The small hammer is used to contact each fingertip perpendicular to the coronal hand plane (Fig.2). In preliminary experiments, hammer impacts are tested to determine if the IMUs’ full scale is reached. Then, approximately the same impacts are manually delivered in all experiments. The experiments consist of one hundred impacts on each fingertip. The three acceleration components are recorded from the five IMUs inside the socket during each impact. Then, the same data processing in Section IV-C is carried out to obtain the one-dimensional signal . Thus, four datasets are obtained with five hundred samples each, one hundred for each fingertip. Each sample comprises five one-dimensional signals from the five IMUs.
V-B LSTM Network Architecture and Hyperparameters
@@ V-C Data Analysis
VI Results
VI-A Tactile Feedback Perception Experiment
@@
Based on the questionnaire, it emerged that the subject feels a vibration on the inner socket of the prosthetic device, which is smooth with the SH and more evident with the CH, and can attribute the vibration to a contact zone.
VI-B Tactile Feedback Transmission Experiment
We recorded the acceleration signals from IMUs glued inside the socket and the impact forces of the pendulum, which contacts the prosthetic fingertips. Fig. 8 shows the different impact forces of the sensorized pendulum with the index finger of each bionic hand.
@@ VI-B Tactile Feedback Transmission Experiment
IMUs’ numbers refer to Fig. 5. Since the socket is always the same, transmission differences depend on the type of hand used. The energy associated with socket accelerations is higher when the rigid VariPlus and Cosmetic hands are worn (especially in the metallic fingers of VP) and decreases with the other hands. Indeed, high-frequency signals are better transmitted in less compliant and less articulated hands. Since the IL thumb is stiffer in the direction of impact, its energy results are greater than the other fingers. For the softer and more articulated hands (SH and IL), the energy is much lower for each IMU. In SH, IMUs energy values are also similar between a finger and another. Indeed, it has several dampening elements, such as the wrist, which prevent vibration transmission through the prosthetic device.
The behaviour of the five IMUs in response to contacts is slightly different from each other in all fingers. This can be noticed especially in the more rigid and less articulated fingers.
The Spearman rank coefficient between the mean energy value of each hand and the accuracy of Tactile Feedback Perception Experiment was 0.8. Thus, a strong monotonic association was found.
VI-C Tactile Feedback Recognition Experiment
We trained four LSTM networks to recognize the fingers contacted based on acceleration data recorded inside the socket, one net for each bionic hand. Nets’ hyperparameters were tuned using Bayesian optimization. Table III shows the hyperparameters tuning results, and Table IV shows validation and test accuracy results and mean precision metrics for the prosthetic hand used.
@@ VI-C Tactile Feedback Recognition Experiment
From the confusion matrices, it should be noted that each class was well predicted from all LSTM models. Precision and recall metrics are also balanced for each class. The overall performances decrease for the SoftHand LSTM model.
VII Discussion
@@
The Tactile Feedback Perception Experiment proved that the subject was able to perceive and recognize the contact on all the bionic hands. The subject detected the finger contacted with positive results. Indeed, the accuracy-related to all the bionic hands is much higher than the chance level of 20%. Despite the daily use of the subject’s cosmetic prostheses, better results were achieved with the CH hand and the reference socket. This might be explained by the reference socket used in the experiments being tighter than the user’s one. Also, the material of the reference socket is stiffer than the personal one and could have affected the perception. Thus, vibrations may be better transmitted to the subject. Results prove that the subject learns, if unknowingly, to associate the transmitted vibration patterns to a finger position with better outcomes for rigid and less articulated hands. Indeed, wearing the Cosmetic and the VariPlus hands, accuracy metrics are 58% and 52% with respect to 45% and 37% of the I-Limb and the SoftHand bionic hands. Regarding the index finger recognition compared to the other fingers, the subject achieved impressive results, with higher accuracy for the CH hand (83%) than VP and IL (65% and 72% accuracies, respectively). With the SH, only a 55% accuracy was achieved, slightly higher than the chance level. Considering the discrimination ability to identify a contact on the thumb from an impact on the other fingers, the higher recognition accuracy for the I-Limb (93%) compared to the CH (81%) and the VP(82%) can be explained by considering that the I-limb thumb has a transverse mechanical play. Thus, during the impact, set on the hand sagittal plane, the thumb is more rigid compared to the other fingers and may make it easier for the subject to detect high-frequency vibration associated with that finger compared to the CH and VP. The SoftHand decreases the performance (accuracy of 61%), but still with a result above the chance level. Obviously, the findings are subjective, and further studies are needed to confirm them for subjects with different amputation types and levels. Moreover, different types of prosthetic impacts and impact directions could be further investigated. The experimental materials affected the results obtained, but we successfully highlighted the differences between the various hands by keeping the same socket.
@@
In the Tactile Feedback Transmission Experiment, we have effectively quantified impacts at the socket level and studied the response of each hand. The results quantitatively prove that softer and more articulated hands dampen contacts while the hands with stiff components and less articulated better transmit high-frequency vibrations. Indeed, more intense accelerations were detected at the socket level for the more rigid hands with respect to the softer ones. The findings reflect the Tactile Feedback Perception Experiment results despite being carried out under different conditions. In the presence of IMUs, the subject could not wear the prosthesis, and the vibration transmission behaviour is different with respect to when the arm is present. Nevertheless, hanging the prosthesis enabled us to characterize each robotic hand and understand how vibration is transmitted regardless of the person wearing the prosthesis. Obviously, when comparing the stiffness differences between a human arm and the air, it is evident that the air-prosthesis connection influenced vibration transmission. Thus, the differences in the experimental setup might have contributed to the lower classification accuracies observed in the Tactile Feedback Perception Experiment. We acknowledge this as a limitation of the study. Future research will address this concern by exploring methods to perform measurements while a prosthetic user wears the socket. Nevertheless, the Spearman rank coefficient of 0.8 demonstrates that user perception and measured accelerations inside the socket are monotonically correlated. The hands with lower energy socket vibrations are those with lower accuracy results and vice versa.
@@
In the Tactile Feedback Recognition Experiment, we proved that LSTM models are able to detect the finger contacted from the high-frequency socket acceleration signals with quite a high accuracy for all hands. Thus, the recorded vibratory patterns contain the information necessary to discriminate the impacted finger, even if registered on the inner surface of the socket. The discrimination ability of ANNs is substantially better than the subject’s one in all prosthetic hands. For acceleration dimensional reduction we used the DTF321 algorithm in both Tactile Feedback Recognition Experiment and Tactile Feedback Transmission Experiment. As mentioned, the DTF321 algorithm is considered one of the best 321 approaches for offline processing when perceptual similarity is prioritized [22]. Still, DTF321 is not rotation invariant, and a different orientation of the IMU yields different DFT321 values even if the signals captured by the IMU are identical in the new orientation of the sensor [21, 22]. Thus, future online implementation will consider the Principal Component Analysis (PCA) algorithm to avoid orientation dependencies [22, 25]. Nevertheless, as a result of our findings, the recorded acceleration signals and ANNs have the potential to compensate for the loss of tactile perception given by the softer and articulated hands. Based on the evidence we provided, the subject already knows how to relate those vibrations to the corresponding finger contacted, especially with rigid hands. Although prosthetic hands with soft and compliant components are adaptable, intuitive to use and more natural in the execution of activities of daily living with respect to rigid hands [26, 27], they do reduce vibration transmission, demoting perception. Thus, enhancing socket vibration transmission or replicating the acceleration signals associated with rigid hands could be a possible solution to improve soft prosthesis finger discrimination performance. Future studies will investigate the use of vibrotactile actuators associated with ANNs to intensify natural high-frequency vibration transmission and perception of tactile cues.
The overall results prove that the characteristics of the prosthetic hand affect users’ perception. Even without a feedback system, the subject was able to perceive impacts and discriminate them with an accuracy higher than the chance level. With softer and more articulated hands, however, accuracy, recall, and precision metrics indicate a lower user’s finger discrimination ability. Softer prosthetic hands provide fewer cues to human discrimination of tactile feedback, but artificial enhancement of vibrotactile feedback is feasible with ANNs trained on socket acceleration data. The findings of this study provide a foundation for new prosthetic design and engineering for artificial feedback. By recognizing prosthetic contact cues with ANN with sensors that are not attached to the prosthetic hand, the integration of a haptic feedback device may become - at least theoretically- feasible, independent of the prosthetic hand used, while not compromising the prosthesis’ wearability or usability. As a further implication, actuators that are not directly attached to the skin could be used to intensify socket vibration, thus preventing any discomfort to a user. Finally, by enhancing the natural transmission of tactile cues, the feedback could be more intuitive and easy to interpret with respect to state-of-the-art non-invasive artificial feedback solutions.
VIII Conclusions
@@
This work successfully tested tactile perception and transmission in four prosthetic hands with different mechanical and kinematics characteristics. The results proved that a prosthetic user perceived tactile stimuli with the four bionic hands without a haptic system. Interestingly, finger discrimination performances were all higher than the chance level, with decreased performances for the softer hands. The four bionic hands were characterized by impact forces and high-frequency stimuli transmission through the socket upon impacts with each finger. It has been quantitatively shown that rigid hands promote tactile transmission while more articulated hands are softer and absorb impacts. These features are often appreciated by users, but they do reduce finger discrimination ability. We have shown that ANNs trained on signals from vibration sensors on the socket can recover maximum finger discrimination ability even with softer articulated hands. The results of this work encourage us to design a new intuitive haptic feedback device able to intensify prosthetic natural vibration transmission performances.

### v2

@@ VI Tactile Feedback Recognition Experiment
VI-A Datasets
New experiments are carried out to obtain four datasets, one for each robotic hand. The prosthetic hands with the reference socket are set by the same supporting structure in Section IV and V. The five IMUs are kept on the inner surface of the socket to measure the vibration pattern after each contact. The small hammer is used to contact each fingertip perpendicular to the coronal hand plane (Fig.1). In preliminary experiments, hammer impacts are tested to determine if the IMUs’ full scale is reached. Then, approximately the same impacts are manually delivered in all experiments. The experiments consist of one hundred impacts on each fingertip. The three acceleration components are recorded from the five IMUs inside the socket during each impact. Then, the same data processing in Section V-C is carried out to obtain the one-dimensional signal . Thus, four datasets are obtained with five hundred samples each, one hundred for each fingertip. Each sample comprises five one-dimensional signals from the five IMUs.
VI-B LSTM Network Architecture and Hyperparameters
@@ VI-C Data Analysis
VII Results
VII-A Tactile Feedback Perception Experiment
@@
Based on the findings from the questionnaire (please refer to Table III), the subject perceived vibrations within the inner socket of the prosthesis. Notably, these vibrations were reported to be smoother when using the SH and more discernible when utilizing the CH. Additionally, the subject could attribute the source of these vibrations to a specific contact zone.
[…]
VII-B Tactile Feedback Transmission Experiment
We recorded the acceleration signals from IMUs glued inside the socket and the impact forces of the pendulum, which contacted the prosthetic fingertips. Fig. 8 shows the different impact forces of the sensorized pendulum with the index finger of each bionic hand.
@@ VII-B Tactile Feedback Transmission Experiment
IMUs’ numbers refer to Fig. 5. Since the socket is always the same, transmission differences depend on the type of hand used. The energy associated with socket accelerations is higher when the rigid VariPlus and Cosmetic hands are worn (especially in the metallic fingers of VP) and decreases with the other hands. Indeed, high-frequency signals are better transmitted in more rigid designs. Since the IL thumb is stiffer in the direction of impact, its energy results are greater than the other fingers. For the softer and more articulated hands (SH and IL), the energy is much lower for each IMU. In SH, IMUs energy values are also similar between a finger and another. Indeed, it has several dampening elements, such as the wrist, which prevent vibration transmission through the prosthetic device.
The behaviour of the five IMUs in response to contacts is slightly different from each other in all fingers. This can be noticed especially in the more rigid and less articulated fingers. The Spearman rank coefficient between the mean energy value of each hand and the accuracy of Tactile Feedback Perception Experiment was 0.8. Thus, a strong monotonic association was found.
VII-C Tactile Feedback Recognition Experiment
We trained four LSTM networks to recognize the fingers contacted based on acceleration data recorded inside the socket, one net for each bionic hand. Nets’ hyperparameters were tuned using Bayesian optimization. Table IV shows the hyperparameters tuning results, and Table V shows validation and test accuracy results and mean precision metrics for the prosthetic hand used.
@@ VII-C Tactile Feedback Recognition Experiment
From the confusion matrices, it should be noted that each class was well predicted from all LSTM models. Precision and recall metrics are also balanced for each class. The overall performance decreased for the SoftHand LSTM model.
VIII Discussion
VIII-A Tactile Feedback Perception Experiment
The Tactile Feedback Perception Experiment proved that the subject was able to perceive and recognize the contact on all the bionic hands. The subject detected the finger contacted with positive results. Indeed, the accuracy related to all the bionic hands is much higher than the chance level of 20%. The subject’s personal prosthesis used in the preliminary experiment includes a CH hand, the same as that used in the experiments, along with a cosmetic socket. Despite the subject being used to their cosmetic prosthesis, which they don regularly, better accuracy was observed with the CH hand and the reference socket in the experiment with respect to that measured in the preliminary one. This difference could be due to the tighter and more rigid structure of the reference socket used in the experiments compared to the user’s own cosmetic socket. Thus, vibrations may be better transmitted to the subject.
@@ VIII-A Tactile Feedback Perception Experiment
Results prove that the subject could associate the transmitted vibration patterns to a finger position with better outcomes for rigid and less articulated hands. Future studies will investigate the potential learning curve of this skill. Wearing the Cosmetic and the VariPlus hands, accuracy metrics are 58% and 52% with respect to 45% and 37% of the I-Limb and the SoftHand bionic hands. Regarding the index finger recognition compared to the other fingers, the subject achieved impressive results, with higher accuracy for the CH hand (83%) than VP and IL (65% and 72% accuracies, respectively). With the SH, only a 55% accuracy was achieved, slightly higher than the chance level. Considering the discrimination ability to identify a contact on the thumb from an impact on the other fingers, the higher recognition accuracy for the I-Limb (93%) compared to the CH (81%) and the VP(82%) can be explained by considering that the I-limb thumb has a transverse mechanical play. Thus, during the impact, set on the hand sagittal plane, the thumb is more rigid compared to the other fingers and may make it easier for the subject to detect high-frequency vibration associated with that finger compared to the CH and VP. The SoftHand decreases the performance (accuracy of 61%), but still with a result above the chance level.
The findings from the questionnaire analysis offer valuable insights into the tactile feedback perception within prosthetic devices. The discernible differences in vibration perception between SH and CH could potentially be attributed to various factors, including the material properties of the sockets and the manner in which vibrations are transmitted through the hands. The more discernible perception with CH indicates a more efficient transmission of vibrations, which may result in reduced sensory fatigue for the user. Moreover, the participant was able to consciously localize the source of vibrations to specific contact zones, underscoring the importance of taking this vibration transmission into account for sensory feedback device design in prosthetic devices. This localization capability may facilitate more intuitive interaction with the prosthesis, allowing users to adapt their movements and behaviours accordingly. Overall, these findings highlight the intricate interplay between design factors and user perception in shaping the efficacy and usability of prosthetic devices.
VIII-B Tactile Feedback Transmission Experiment
@@
In the Tactile Feedback Transmission Experiment, we have effectively quantified impacts at the socket level and studied the response of each hand. The results quantitatively reinforce our intuition that advanced, softer, and more articulated hands dampen contacts, while rigid hands with stiff components and fewer articulations better transmit high-frequency vibrations. Indeed, more intense accelerations were detected at the socket level for the more rigid hands with respect to the advanced ones. While the observed outcome aligns with expectations, it’s crucial to note the lack of prior empirical evidence directly supporting this hypothesis. The findings reflect the Tactile Feedback Perception Experiment results despite being carried out under different setup conditions. In the presence of IMUs, the subject could not wear the prosthesis, and the vibration transmission behaviour is different with respect to when the arm is present. Nevertheless, hanging the prosthesis enabled us to characterize each robotic hand and understand how vibration is transmitted regardless of the person wearing the prosthesis. The Spearman rank coefficient of 0.8 demonstrates that user perception and measured accelerations inside the socket are monotonically correlated. The hands with lower energy socket vibrations are those with lower accuracy results and vice versa.
VIII-C Tactile Feedback Recognition Experiment
In the Tactile Feedback Recognition Experiment, we proved that LSTM models are able to detect the finger contacted from the high-frequency socket acceleration signals with quite a high accuracy for all hands. Thus, the recorded vibratory patterns contain the information necessary to discriminate the impacted finger, even if registered on the inner surface of the socket. The discrimination ability of ANNs is substantially better than the subject’s one in all prosthetic hands. The distinct data collection methods in this experiment compared to the Tactile Feedback Transmission Experiment are driven by their divergent objectives. In the Tactile Feedback Transmission Experiment, the use of the sensorized pendulum facilitated consistent quantification of each impact, allowing for analysis and comparison across different hands. Conversely, in the Tactile Feedback Recognition Experiment, we aimed to replicate the conditions of the Tactile Feedback Perception Experiment, necessitating the introduction of differences between impacts using a small hammer rather than producing identical impacts.
@@ VIII-C Tactile Feedback Recognition Experiment
For acceleration dimensional reduction, we used the DTF321 algorithm in both Tactile Feedback Recognition Experiment and Tactile Feedback Transmission Experiment. As mentioned, the DTF321 algorithm is considered one of the best 321 approaches for offline processing when perceptual similarity is prioritized [24]. Still, DTF321 is not rotation invariant, and a different orientation of the IMU yields different DFT321 values even if the signals captured by the IMU are identical in the new orientation of the sensor [20, 24]. Thus, future online implementation will consider the Principal Component Analysis (PCA) algorithm to avoid orientation dependencies [24, 25]. Nevertheless, as a result of our findings, the recorded acceleration signals and ANNs have the potential to compensate for the loss of tactile perception given by the advanced hands.
VIII-D Limitations
About the Tactile Feedback Perception Experiment, we acknowledge the limitation arising from the inclusion of only one prosthetic user in the experiment. Obviously, the findings are subjective, emphasizing the necessity for future studies to validate them with an increased and heterogeneous group of participants, encompassing individuals with varying nature, levels, location of amputations and state of the nerves in the residual limb, which can affect the user’s perception. Moreover, it is important to consider that the participant’s perception may be influenced by external variables, such as their state of mind, discomfort, or boredom, alongside factors related to skin condition, thus adding complexity to the interpretation of the results. Furthermore, it is essential to acknowledge that our assessment of prosthetic user performance was conducted under controlled conditions, specifically in a quiet room, where the participant was fully engaged in the designated task. This deliberate choice was made to minimize extraneous factors that could potentially detract from the accuracy of our findings, although at the expense of mirroring real-world situations. Our future research will encompass experiments conducted in environments that more accurately replicate real-life scenarios.
Deliberately choosing a small hammer over the sensorized pendulum used in the Tactile Feedback Transmission Experiment aimed to avoid significant delays in setup and repositioning for each contact. This decision was made to maintain the subject concentration and minimize fatigue during the experimental session. Additionally, it allowed for replicating realistic impact scenarios like those encountered in everyday environments and suggested that the ANNs are also able to discriminate among different levels of impact. However, experimenter errors may have prevented consistent impact levels and directions for each finger and repetition. Future research will explore alternative methods to deliver contact in a more quantified manner while still prioritizing the avoidance of fatigue or discomfort for the subject. Furthermore, our research will explore various types of prosthetic impacts and directions. Exploring changes in impact location, orientation, amplitude, and contact size on the prosthetic hand could yield valuable insights in future studies. We also aim to examine how simultaneous impacts interact to produce sustained and continuously modulated cues and whether similar identifiable propagation patterns emerge during shear loading.
The choice of experimental materials, notably the type and material of the socket, influenced our findings. It is reasonable to posit that variations in these factors may have significantly impacted the transmission of vibrations through the socket, thereby shaping the perceptual responses of the subjects involved in our experiments. This assertion is substantiated by our study, which demonstrated the influence of different prosthetic hands on perception. Future work will be focused on conducting additional tests aimed at systematically investigating the effects of diverse socket types on both vibration transmission and subject perception. This will necessitate a comprehensive exploration of various socket designs and materials to gain deeper insights into their influence on the overall performance of prosthetic devices. Nevertheless, by consistently employing the same reference socket throughout our experiments, we were able to effectively discern differences among the various prosthetic hands used.
The presence of inertial measurement units (IMUs) prevented the subject from wearing the prosthesis, potentially leading to improved classification accuracies in the Tactile Feedback Recognition Experiment. However, this setup simplified post-impact oscillation filtering by using a high-pass filter to remove low-frequency socket oscillations. Indeed, the oscillations from the suspended socket differ greatly in frequency from the high-frequency oscillations associated with the vibration modes of the prosthesis. Any link to a rigid support, like the human arm, would accelerate the damping of the high-frequency oscillations via contact transmission, complicating measurement. However, it’s crucial to note that these transmitted vibrations are precisely what elicit human perception, aligning with our measurement objective. Addressing the potential alteration of prosthesis dynamics in the absence of a human arm, future research will explore methods for measuring impulse propagation while a prosthetic user wears the socket, ensuring comfort is not compromised.
VIII-E Study Significance and Future Directions
Our results indicate that part of the tactile information in exploring the external world is transmitted through the prosthesis and is not lost, with differences based on the type of hand. In rigid prosthetic designs, the transmitted information has allowed the subject to associate vibration transmission with specific tactile cues. Consequently, this may diminish the necessity of haptic feedback devices (e.g. vibrotactile actuators for transmitting initial contact or texture cues). However, in the case of softer and more articulated prosthetic designs, this correlation is not consistently observed. Although prosthetic hands with soft and compliant components are adaptable, intuitive to use and more natural in the execution of activities of daily living with respect to rigid hands [26, 27], they do reduce vibration transmission, demoting perception. Therefore, for users utilizing advanced prosthetic designs or those who still desire tactile feedback, it’s crucial to recognize that feedback device designs should consider the natural vibration transmission through the prosthesis. Although less intense than in rigid designs, this transmission remains persistent and offers potential for enhancement, enriching user experience without adding artificial signals unrelated to natural vibration transmission. One potential approach may involve enhancing natural vibration transmission by integrating vibrotactile actuators within the socket controlled by accelerometer data and adjustable gain settings. Furthermore, we speculate that training a neural network to identify impacts based on socket accelerations from various prosthetic hands could lead to the development of a device that is independent of the specific type of prosthesis used, offering a more adaptable and resilient solution. Nevertheless, we acknowledge the difficulty of deciding which hand propagation pattern to prioritize. Overall, these potential solutions serve the same purpose of utilizing existing vibrations and potentially decreasing users’ cognitive load required to interpret artificial tactile information transmitted by tactile feedback devices. Further research is necessary to fully comprehend the phenomenon and effectively utilize the results obtained, and we intend to address these challenges in future investigations. Additionally, results show that detecting and recognizing tactile cues from the accelerometers at the bottom of the socket is feasible. This indicates that valuable data can be extracted from a region between the prosthetic hand and the end of the socket. Considering that haptic devices for prosthetics currently on the market have sensors placed on the prosthetic hand fingers, we believe that placing them on the socket could enhance the robustness of the prosthesis, reducing the risk of breakage or malfunctions, and increase the wearability and integration of this type of devices.
IX Conclusions
@@
This work successfully tested tactile perception and transmission in four prosthetic hands with different characteristics. The results proved that a prosthetic user perceived tactile stimuli with the four bionic hands without a haptic system. Interestingly, finger discrimination performances were all higher than the chance level, with decreased performances for the advanced hands. The four bionic hands were characterized by impact forces and high-frequency stimuli transmission through the socket upon impacts with each finger. It has been quantitatively shown that rigid hands promote tactile transmission while more advanced hands absorb impacts. We have shown that ANNs trained on signals from vibration sensors on the socket can recover maximum finger discrimination ability even with softer articulated hands. This indicates that rigid prosthetic designs favour natural vibration transmission, potentially reducing the need for extra tactile feedback devices, while advanced designs may lack this phenomenon, presenting an opportunity to enhance weaker tactile transmission and create intuitive haptic devices that minimize adaptation to artificial signals.